Find the effective rate corresponding to the given nominal rate. (Round your answer to two decimal places.)
11%/year compounded semiannually
Answer #1
I'm so lost...
answered by: Marie20
Answer #2
make a comparison base on one year
let the effective rate be i

(1+i)^1 = (1.055^2
1+i = 1.113025
i = .113025 or 11.3025% per annum
answered by: tob